OpenStudy (anonymous):

Use substitution to solve the system of equations. Enter your answer as an ordered pair; example: (4, -5) y = x - 19 2x - y = 27

OpenStudy (texaschic101):

y = x - 19 2x - y = 27 sub in x - 19 in for y in the second equation 2x - (x - 19) = 27 2x - x + 19 = 27 x = 27 - 19 x = 8 now sub in 8 for x in either of the original equations y = x - 19 y = 8 - 19 y = -11 check.. 2x - y = 27 2(8) - (-11) = 27 16 + 11 = 27 27 = 27 (correct) answer is (8,-11)